### 14 Mar 2026 v1

# SHAMISA: SHAped Modeling of Implicit Structural Associations for Self-supervised No-Reference Image Quality Assessment

Abstract
No-Reference Image Quality Assessment (NR-IQA) aims to estimate perceptual quality without access to a reference image of pristine quality. Learning an NR-IQA model faces a fundamental bottleneck: its need for a large number of costly human perceptual labels. We propose SHAMISA, a non-contrastive self-supervised framework that learns from unlabeled distorted images by leveraging explicitly structured relational supervision. Unlike prior methods that impose rigid, binary similarity constraints, SHAMISA introduces implicit structural associations, defined as soft, controllable relations that are both distortion-aware and content-sensitive, inferred from synthetic metadata and intrinsic feature structure. A key innovation is our compositional distortion engine, which generates an uncountable family of degradations from continuous parameter spaces, grouped so that only one distortion factor varies at a time. This enables fine-grained control over representational similarity during training: images with shared distortion patterns are pulled together in the embedding space, while severity variations produce structured, predictable shifts. We integrate these insights via dual-source relation graphs that encode both known degradation profiles and emergent structural affinities to guide the learning process throughout training. A convolutional encoder is trained under this supervision and then frozen for inference, with quality prediction performed by a linear regressor on its features. Extensive experiments on synthetic, authentic, and cross-dataset NR-IQA benchmarks demonstrate that SHAMISA achieves strong overall performance with improved cross-dataset generalization and robustness, all without human quality annotations or contrastive losses.
I Introduction
Image Quality Assessment (IQA) aims to estimate perceptual image quality in line with human opinion. The No-Reference setting (NR-IQA) is especially challenging, as it must operate without reference images or distortion labels. NR-IQA is critical for real-world tasks such as perceptual enhancement [1], image captioning [2], and streaming optimization [3]. However, modeling perceptual quality is difficult due to the complex interplay between distortion types and content [4, 5]. Supervised approaches [6, 7, 8] rely on extensive human annotations, with KADID-10K [9] alone requiring over 300,000 subjective ratings, making them expensive and hard to scale. To address this, recent efforts turn to self-supervised learning (SSL), using unlabeled distorted images to learn quality-aware representations [10, 11], typically via contrastive objectives tailored to either content or degradation similarity.
Classical SSL methods like SimCLR [12] and MoCo [13], developed for classification, learn content-centric and distortion-invariant features. Such representations misalign with NR-IQA, where both content and degradation must be modeled. To bridge this gap, contrastive SSL-IQA methods introduce domain-specific objectives but can suffer from sampling bias because sampled negatives may include semantically related false negatives [14, 15, 16]. CONTRIQUE [17] groups images by distortion type and severity, ignoring content. Re-IQA [11], by contrast, uses overlapping crops of the same image to promote content-aware alignment. While each mode captures useful relations, similarity is enforced either across content or across distortions, but not both. This leads to scattered embeddings for similarly degraded images with different content [18]. ARNIQA [18] learns a distortion manifold by aligning representations of similarly degraded images, irrespective of content. Yet, such approaches rigidly collapse embeddings without accounting for perceptual effects like masking [4], where content alters perceived quality. In practice, perceptual similarity depends jointly on both distortion severity and image content, a dependency that current SSL-IQA models fail to capture in a flexible, scalable manner.
We propose SHAMISA (SHAped Modeling of Implicit Structural Associations), a non-contrastive self-supervised framework that addresses this challenge by learning representations jointly sensitive to both distortion and content through explicitly constructed relation graphs. “SHAped Modeling” refers to graph-based relational supervision, whereas “Implicit Structural Associations” denote the latent perceptual relations encoded in these graphs, providing finer control over similarity learning than prior SSL-IQA methods that treat distorted views uniformly. SHAMISA draws inspiration from ExGRG [19], which introduced explicit graph-based guidance for self-supervised learning in the graph domain. In contrast, SHAMISA extends this idea to visual data, where quality prediction depends on fine-grained perceptual variations. Prior methods implicitly impose sparse relational structures, resulting in disconnected or inconsistent supervision, which SHAMISA addresses through explicitly shaped relation graphs.
At the core of SHAMISA is a compositional distortion engine that generates infinite compositional degradations from continuous parameter spaces. Each mini-batch is built from reference images that form distortion composition groups where only one degradation factor varies, enabling controlled sampling across content, distortion type, and severity. From this, we build two categories of relation graphs: (i) Metadata-Driven Graphs, which encode pairwise similarity based on distortion metadata, encouraging images with similar degradations to lie close in the learned manifold while inducing controlled representational shifts as distortion severity varies; and (ii) Structurally Intrinsic Graphs, constructed from the latent feature space using $k$-nearest neighbors (kNN) and deep clustering. These graphs supervise a non-contrastive VICReg-style objective, optimized in a single update that constructs relation graphs from the current representations and applies stop-gradient through graph construction.
SHAMISA unifies content-dependent and distortion-dependent learning within a single relational framework, generalizing the rigid pairing schemes used in prior SSL-IQA methods. At test time, the learned encoder is frozen and paired with a linear regressor to predict quality scores. SHAMISA achieves strong performance across both synthetic and authentic datasets without requiring human quality labels or contrastive objectives.
Our main contributions are:
  1. 1.
    We introduce SHAMISA, a non-contrastive self-supervised framework for NR-IQA that encodes both distortion-aware and content-aware information into a shared representation space via explicit relational supervision.
  2. 2.
    We propose a distortion engine that generates compositional groups with controlled variation, enabling fine-grained similarity learning across distortion factors and content.
  3. 3.
    We develop a dual-source relation graph construction strategy that combines Metadata-Driven Graphs based on known degradation metadata and Structurally Intrinsic Graphs derived from the evolving feature space, trained with a stop-gradient alternating update that couples on-the-fly graph construction with representation learning.
  4. 4.
    SHAMISA generalizes the rigid pairing schemes used in prior SSL-IQA methods within a unified graph-weighted invariance framework and achieves strong overall performance across synthetic, authentic, and cross-dataset benchmarks, including the strongest overall six-dataset average among the compared SSL methods, together with improved robustness and transfer.
II Related Work
II-A Traditional and Supervised NR-IQA
No-Reference Image Quality Assessment (NR-IQA) aims to estimate perceptual image quality without access to pristine references. Traditional methods rely on handcrafted features derived from Natural Scene Statistics (NSS), modeling distortions as deviations from expected regularities. Examples include BRISQUE [20], NIQE [21], DIIVINE [22], and BLIINDS [23], while CORNIA [24] and HOSA [25] construct codebooks from local patches. Though effective on synthetic distortions, these models often fail on authentically distorted images due to a lack of semantic awareness.
Supervised NR-IQA models typically use pre-trained CNNs (e.g., ResNet [26]) to extract deep features, which are mapped to quality scores via regression. Techniques like HyperIQA [7], RAPIQUE [27], and MUSIQ [28] adapt neural architectures for perceptual modeling. PaQ-2-PiQ [29] incorporates both image- and patch-level quality labels. Others like DB-CNN [6] and PQR [30] combine multiple feature streams. However, such methods require large-scale and costly human annotations, limiting scalability and generalization.
II-B Self-Supervised Learning for NR-IQA
Self-supervised learning (SSL) enables representation learning from unlabeled data. Contrastive methods like SimCLR [12] and MoCo [13] rely on sampled negatives, which can introduce false negatives and sampling bias when semantically related samples are treated as negatives [14, 15, 16]. Non-contrastive methods like VICReg [31] avoid negatives through invariance and decorrelation. In graph domains, ExGRG [19] introduces explicit relation graphs for SSL, combining domain priors and feature structure. SHAMISA adopts a similar inductive bias but applies it to NR-IQA, where perceptual similarity depends jointly on content and degradation.
Recent SSL-IQA works adapt these paradigms to quality prediction. CONTRIQUE [17] clusters samples with similar distortion types into classes, learning quality-aware embeddings. QPT [10] aligns patches under shared distortion assumptions. Re-IQA [11] uses two encoders: one pre-trained for content-specific features and another trained for distortion-specific features. Their outputs are concatenated and fed to a single linear regressor. This assumes that a shallow fusion layer can recover complex content-distortion interactions from separately learned streams. Re-IQA may therefore under-represent content-distortion interactions, even though these interactions play a central role in perceptual quality assessment rather than being evaluated in isolation [32, 33].
ARNIQA [18] instead aligns representations of images degraded equally, disregarding content. While it captures distortion similarity, it rigidly enforces uniform proximity and ignores perceptual effects introduced by content. In contrast, SHAMISA learns a unified representation space that respects both distortion and semantic structure, using two complementary relation graphs. A metadata-driven graph and a structurally intrinsic graph together encode distortion-aware similarity and perceptual affinities across content. This provides soft, fine-grained constraints that generalize prior methods as special cases.
II-C Compositional Distortion Modeling and Relational Supervision
Degradation engines are central to SSL-IQA training. Prior models apply a limited set of discrete distortions or fixed composition rules [11, 10]. ARNIQA, for example, applies sequential distortions with fixed sampling schemes but lacks structured sampling for controlled variation. RealESRGAN [1] follows a fixed distortion sequence from predefined groups. Such engines offer limited degradation diversity and do not support relational supervision.
SHAMISA introduces a compositional distortion engine that generates continuously parameterized degradations with uncountable variation. We partition each mini-batch into tiny-batches (small, fixed-size subsets). Each tiny-batch contains distortion composition groups in which only one degradation factor varies, enabling precise control over severity and type. This supports precise metadata generation for graph construction. Unlike prior works that enforce binary or fixed similarity constraints, SHAMISA enables structured shifts in representation space by applying soft supervision from graph relations. Similar samples lie closer, while severity variation introduces gradual transitions.
Combined with its dual-graph relational supervision, SHAMISA learns distortion-aware, content-sensitive embeddings in a non-contrastive setting. By decoupling quality learning from rigid class labels and negative sampling, SHAMISA offers a scalable and generalizable framework for NR-IQA without relying on human quality annotations or conventional augmentation heuristics.
III Method
Figure 1: Overview of the proposed SHAMISA framework. (1) Self-supervised pre-training: pristine images are transformed by a compositional distortion engine to form a mini-batch $\mathbf{X}$ and distortion metadata; the encoder $f_{\theta}$ and projector $g_{\psi}$ produce representations $\mathbf{H}$ and embeddings $\mathbf{Z}$ used to construct dual-source relation graphs (metadata-driven and structurally intrinsic), which are aggregated by a multi-source aggregator $\Phi$ into $\mathbf{G}$ and optimize a graph-weighted SSL objective. (2) Downstream NR-IQA: we freeze $f_{\theta}$ and train a lightweight regressor on top of $\mathbf{H}$ to predict quality scores.
As illustrated in Fig. 1, SHAMISA combines structured distortion generation with dual-source relational supervision to learn quality-aware features in a non-contrastive setting. We then transfer the learned encoder to NR-IQA by freezing backbone representations and fitting a lightweight regressor, so downstream performance reflects representation quality rather than head complexity.
III-A Overview and SSL Formulation for NR-IQA
We pre-train a ResNet-50 encoder $f_{\theta}$ with a 2-layer MLP projector $g_{\psi}$ on unlabeled images degraded online by our distortion engine, optimizing a VICReg-style non-contrastive objective [31] to avoid negative sampling and its sampling-bias issues [14]. After this self-supervised pre-training, we discard $g_{\psi}$, freeze $f_{\theta}$ (never fine-tune), and train a linear regressor on the frozen features for quality prediction.
Tensorized batch construction
Let $N_{\text{ref}}$ denote the number of reference images per mini-batch and $N_{\text{comp}}$ the number of distortion compositions applied to each reference image (Sec. III-B). We take one random crop per pristine image to form
${\bm{R}}^{s}\in\mathbb{R}^{N_{\text{ref}}\times 3\times H_{in}\times W_{in}}.$
Sampling $N_{\text{comp}}$ compositions $\{\chi_{d}\}_{d=1}^{N_{\text{comp}}}$ yields distorted sets
${\bm{X}}^{(d)}\,=\,\chi_{d}({\bm{R}}^{s})\in\mathbb{R}^{N_{\text{ref}}\times 3\times H_{in}\times W_{in}},\quad d=1,\dots,N_{\text{comp}}.$
We concatenate references and all distorted sets:
${\bm{X}}$ $=[\,{\bm{R}}^{s};\ {\bm{X}}^{(1)};\ \dots;\ {\bm{X}}^{(N_{\text{comp}})}\,]\in\mathbb{R}^{N\times 3\times H_{in}\times W_{in}},$ (1)
$N$ $=N_{\text{ref}}\,(N_{\text{comp}}+1).$
Representations $\bm{H}$} and Embeddings $\bm{Z}$} are
${\bm{H}}=f_{\theta}({\bm{X}})\in\mathbb{R}^{N\times d_{h}},\qquad{\bm{Z}}=g_{\psi}({\bm{H}})\in\mathbb{R}^{N\times d_{z}}.$
Human opinion scores are used only for the regression head after pre-training.
III-B Compositional Distortion Engine
A distortion function is one atomic degradation from KADID-10K with 24 functions across 7 categories (Brightness change, Blur, Spatial, Noise, Color, Compression, Sharpness & Contrast) [9]. A distortion composition in SHAMISA is an ordered composition of multiple distortion functions with at most one function per category, applied sequentially. Unlike prior SSL-IQA setups based on discrete severity grids [10, 18], SHAMISA samples continuous severities, yielding an uncountable family of compositions. Order is perceptually consequential, hence we randomize it. In practice each iteration uses a finite $N_{\text{comp}}$, but randomization across iterations explores novel compositions, and continuous level differences are exploited later by our relation graphs.
Severity calibration
For each distortion function $f_{m}$ we map its native parameter domain $\lambda_{m}^{\min},\,\lambda_{m}^{\max}]$ into $[0,1]$ via a piecewise-linear calibration $\rho_{m}:[\lambda_{m}^{\min},\,\lambda_{m}^{\max}]\to[0,1]$ obtained by linearly interpolating the discrete intensities used in [9]. This yields comparable per-function normalized severities. During sampling we operate in the normalized space; if needed, we recover native parameters by $\rho_{m}^{-1}$.
Sampling a composition
For each distortion composition we sample a tuple $(M,\,S,\,f,\,\pi,\,\lambda)$ where: $M\sim\mathrm{Unif}\{1,\dots,M_{d}\}$ is the number of distortion functions in the composition with at most one per category; $S\subset\{1,\dots,7\}$ is a size-$M$ subset indexing the seven KADID-10K distortion categories, sampled uniformly without replacement; $f=(f_{1},\dots,f_{M})$ picks one function per category in $S$; $\pi$ is a permutation of $\{1,\dots,M\}$ sampled uniformly indicating the application order; and $\lambda=(\lambda_{1},\dots,\lambda_{M})\in[0,1]^{M}$ are per-function normalized severities sampled i.i.d. by drawing $\epsilon_{m}\sim\mathcal{N}(0,0.5^{2})$ and setting $\lambda_{m}=\min(1,|\epsilon_{m})$, where $0,1]^{M}$ is the $M$-dimensional unit hypercube. Given an input image $x$, let $\theta_{m}\coloneqq\rho_{m}^{-1}(\lambda_{m})$ and write $f_{m}(\,\cdot\,;\theta_{m})$. The composed degradation is
$\mathcal{C}(x;\pi,f,\lambda)\;=\;\big(f_{\pi(M)}(\,\cdot\,;\theta_{\pi(M)})\circ\cdots\circ f_{\pi(1)}(\,\cdot\,;\theta_{\pi(1)})\big)(x).$ (2)
III-B1 Single-factor variation and trajectories
Each mini-batch is partitioned into $B$ tiny-batches; we use $n]\coloneqq\{1,\dots,n\}$ for index ranges, so $i\in[B]$ indexes tiny-batches. In tiny-batch $i$, select $R$ references $\{r_{i,j}\}_{j=1}^{R}$ and instantiate $C$ distortion composition groups. For the $k$-th composition group ($k\in[C]$), sample a base composition $(M_{i,k},S_{i,k},f_{i,k},\pi_{i,k},\lambda_{i,k}^{(0)})$, where $\lambda_{i,k}^{(0)}\in[0,1]^{M_{i,k}}$ is the base severity vector. Choose exactly one varying coordinate $m^{\star}(i,k)\in\{1,\dots,M_{i,k}\}$, that is, exactly one distortion function whose level will vary within the composition group, and generate $L$ severity levels $\{s_{i,k}^{(l)}\}_{l=1}^{L}\subset[0,1]$. We draw each level i.i.d. as $s_{i,k}^{(l)}=\min(1,|\epsilon^{(l)}|)$ with $\epsilon^{(l)}\sim\mathcal{N}(0,0.5^{2})$. Define the per-level severities component-wise by
$(\lambda_{i,k}^{(l)})_{m}\;=\;\begin{cases}(\lambda_{i,k}^{(0)})_{m},&m\neq m^{\star}(i,k),\\[2.0pt] s_{i,k}^{(l)},&m=m^{\star}(i,k).\end{cases}$
This single-factor scheme varies one distortion function while holding the others fixed, producing controlled severity variation that isolates that function’s effect in the learned representation. Formally, applying the level-$l$ composition in group $k$ to reference $r_{i,j}$ yields the distorted image
$x_{i,j,k,l}$ $=\mathcal{C}\big(r_{i,j};\,\pi_{i,k},f_{i,k},\lambda_{i,k}^{(l)}\big),$ (3)
$\qquad i\!\in\![B],\ j\!\in\![R],\ k\!\in\![C],\ l\!\in\![L].$
Consequently, the mini-batch contains $N_{\text{ref}}=BR$ references and each reference receives $N_{\text{comp}}=CL$ compositions, with sampling performed independently for each tiny-batch.
III-C Explicit Relation Graphs and Graph-weighted VICReg
We adopt the variance and covariance regularizers of VICReg [31] and replace its augmentation-paired invariance term $\mathcal{L}_{\mathrm{inv}}$ with a graph-weighted variant $\mathcal{L}^{\prime}_{\mathrm{inv}}$, in the spirit of explicitly generated relation graphs [19]:
$\mathcal{L}_{\mathrm{var}}$ $=\sum_{t=1}^{d_{z}}\max\big(0,\,1-\sqrt{\mathrm{Cov}({\bm{Z}})_{t,t}}\big),$ (4)
$\mathcal{L}_{\mathrm{cov}}$ $=\sum_{t=1}^{d_{z}}\sum_{\begin{subarray}{c}u=1\\ u\neq t\end{subarray}}^{d_{z}}\mathrm{Cov}({\bm{Z}})^{2}_{t,u},$ (5)
$\mathcal{L}_{\mathrm{inv}}$ $=\sum_{(i,j)\in\mathcal{A}}\big\|{\bm{Z}}_{i,.}-{\bm{Z}}_{j,.}\big\|_{2}^{2},$ (6)
$\mathcal{L}^{\prime}_{\mathrm{inv}}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}{\bm{G}}_{i,j}\,\big\|{\bm{Z}}_{i,.}-{\bm{Z}}_{j,.}\big\|_{2}^{2}.$ (7)
Here $\mathcal{A}\subseteq[N]\times[N]$ denotes the set of augmentation positives, and ${\bm{G}}\in[0,1]^{N\times N}$ is the weighted adjacency matrix of a soft relation graph (Sec. III-F), where each entry specifies how strongly the corresponding pair is encouraged to remain invariant. These soft, controllable entries of $\bm{G}$} instantiate the implicit structural associations, encoding relational cues that jointly capture content and distortion. Intuitively, $\mathcal{L}_{\mathrm{var}}$ enforces a minimum per-dimension variance to avoid collapse, $\mathcal{L}_{\mathrm{cov}}$ decorrelates dimensions, and $\mathcal{L}^{\prime}_{\mathrm{inv}}$ brings together embeddings proportionally to relational strengths encoded by $\bm{G}$}. In contrast, $\mathcal{L}_{\mathrm{inv}}$ uses a rigid binary set $\mathcal{A}$ of augmentation positives. When $\bm{G}$} is the binary indicator of $\mathcal{A}$ (i.e., ${\bm{G}}_{i,j}=1$ iff $(i,j)\in\mathcal{A}$ and 0 otherwise), Eq. 7 reduces to VICReg’s original invariance in Eq. 6.
Our self-supervised pre-training objective is
$\mathcal{L}_{\mathrm{ssl}}\;=\;\alpha\,\mathcal{L}_{\mathrm{var}}+\beta\,\mathcal{L}_{\mathrm{cov}}+\gamma\,\mathcal{L}^{\prime}_{\mathrm{inv}}.$ (8)
SHAMISA remains non-contrastive: we do not use negatives and instead prevent collapse via the variance-covariance terms [31], rather than through negative pairs as in contrastive methods [12].
Relation to prior positives
Cosine invariance underlies InfoNCE in SimCLR [12]. With $\ell_{2}$-normalized features, $\|a-b\|_{2}^{2}=2\,(1-\cos(a,b))$, so Euclidean and cosine invariances are equivalent up to scaling. Thus, prior invariances that rely on rigid binary pairings become special cases of SHAMISA’s soft graph-weighted invariance. ARNIQA pairs images with the same distortion type and level across different contents [18]; Re-IQA pairs overlapping crops of the same image (content-driven) [11]; CONTRIQUE uses synthetic distortion classes or instance discrimination for UGC [17, 10]. Each choice is recovered by setting $\bm{G}$} to the corresponding sparse binary adjacency.
Why an explicit $\bm{G}$}?
Laplacian Eigenmap view: In two-view augmentation setups, the augmentation graph splits into exactly $N/2$ disconnected components, so its Laplacian has $N/2$ zero eigenvalues [34]. Allowing soft cross-component edges in $\bm{G}$} alleviates this rank deficiency and yields a better posed invariance term [19]. Alternating update view: We construct $\bm{G}$} from the current representations to define soft relational targets, then update $f_{\theta}$ and $g_{\psi}$ by minimizing Eq. 8. A stop-gradient on the graph inputs, defined formally in Sec. III-F, lets graph construction and parameter updates be executed in a single optimization step [35, 19].
III-D Metadata-Driven Graphs
We convert engine metadata into soft relational weights that shape the embedding space through two complementary effects. Within a given content, similarity to the pristine anchor decreases smoothly as severity increases. Across contents, samples from the same distortion composition group remain neighbors when their severities are close.
We next define two monotone maps $\phi,\hat{\phi}:[0,1]\to[0,1]$, with $\phi^{\prime}(u)\leq 0$ and $\hat{\phi}^{\prime}(u)\leq 0$, that convert a normalized severity or severity gap into a similarity weight. For simplicity, we set $\phi(u)=\hat{\phi}(u)=\exp(-\kappa u)$; both maps are bounded in $[0,1]$ and strictly decreasing in $u$.
III-D1 Reference-distorted graph ${\bm{G}}^{rd}$
Recall that in tiny-batch $i$ and group $k$, $m^{\star}(i,k)$ is the varying function and $\lambda^{(l)}_{i,k}\in[0,1]^{M_{i,k}}$ is the level-$l$ severity vector. We assign an edge from each reference to each of its degraded versions:
${\bm{G}}^{rd}\big[r_{i,j},\,x_{i,j,k,l}\big]\;=\;\phi\Big((\lambda^{(l)}_{i,k})_{m^{\star}(i,k)}\Big).$ (9)
Intuition. Small severities yield $\phi\!\approx\!1$ and keep mild degradations near the pristine anchor; large severities drive $\phi\!\to\!0$, allowing strong degradations to move away. This produces smooth, severity-aware attraction without collapse.
III-D2 Distorted-distorted graph ${\bm{G}}^{dd}$
Within the same distortion composition group $(i,k)$, define the 1D severity gap
$\delta_{i,k}(l,l^{\prime})\;=\;\big|\,s^{(l)}_{i,k}-s^{(l^{\prime})}_{i,k}\,\big|\in[0,1],$
and, for any $j,j^{\prime}\in[R]$, define the edge weight between two distorted images whose contents may differ as
${\bm{G}}^{dd}\big[x_{i,j,k,l},\,x_{i,j^{\prime},k,l^{\prime}}\big]\;=\;\hat{\phi}\Big(\delta_{i,k}(l,l^{\prime})\Big).$ (10)
Intuition. Nearby severity levels attract; distant levels do not. This ties together similarly degraded images regardless of content, complementing ${\bm{G}}^{rd}$. To avoid density, we apply a global Top-$K_{d}$ sparsifier:
${\bm{G}}^{dd}\;\leftarrow\;F_{K_{d}}({\bm{G}}^{dd}),$
where $F_{K_{d}}$ retains the $K_{d}$ largest entries in the whole matrix and zeros the rest.
III-D3 Reference-reference graph ${\bm{G}}^{rr}$
We weakly connect pristine images across contents to stabilize a common high-quality anchor:
${\bm{G}}^{rr}\big[r_{i,j},\,r_{i^{\prime},j^{\prime}}\big]\;=\;w_{rr},\qquad w_{rr}\in(0,1].$ (11)
Intuition. This graph builds a coherent pristine neighborhood across contents without forcing content collapse; variance-covariance regularizers prevent trivial solutions while these links stabilize a shared high-quality anchor in the representation space.
The three metadata graphs provide severity-aware, content-aware soft relations that are fed to the overall relation aggregation (Sec. III-F) and used inside the graph-weighted invariance term (Eq. 7) to shape quality-sensitive embeddings.
III-E Structurally Intrinsic Graphs
Beyond metadata, we inject emergent structure from the evolving representation space. These graphs encourage local manifold smoothness and global prototype organization, complementing the metadata graphs.
III-E1 kNN relation graph
Compute cosine similarities on encoder features $\bm{H}$}:
${\bm{S}}^{H}_{i,j}\,=\,\frac{{\bm{H}}_{i,.}^{\top}{\bm{H}}_{j,.}}{\|{\bm{H}}_{i,.}\|_{2}\,\|{\bm{H}}_{j,.}\|_{2}}.$
Form the kNN graph by keeping, for each sample, the top $k_{n}$ similarity scores and removing self-connections:
${\bm{G}}^{k}\,=\,F_{k_{n}}({\bm{S}}^{H})\in\mathbb{R}^{N\times N},$ (12)
where $F_{k_{n}}$ retains the largest $k_{n}$ entries in each row of ${\bm{S}}^{H}$ and zeros the rest.
Intuition. Deep features are well correlated with perceptual similarity [36]; thus nearest neighbors in $\bm{H}$} define perceptually coherent local neighborhoods. ${\bm{G}}^{k}$ promotes local smoothness of the embedding with respect to this current geometry.
III-E2 OT-based clustering guidance
Following ExGRG [19], we adapt optimal-transport (OT) clustering guidance to IQA. Let ${\bm{C}}\in\mathbb{R}^{K\times d_{h}}$ be $K$ trainable prototypes and ${\bm{A}}\in\mathbb{R}^{N\times K}$ the soft assignments
${\bm{A}}_{j,k}\,=\,\frac{\exp({\bm{H}}_{j,.}^{\top}{\bm{C}}_{k,.}/\tau_{c})}{\sum_{k^{\prime}}\exp({\bm{H}}_{j,.}^{\top}{\bm{C}}_{k^{\prime},.}/\tau_{c})}.$ (13)
Balanced targets ${\bm{T}}\in\mathbb{R}^{N\times K}$ are obtained by a Sinkhorn-Knopp solver with entropy regularization [37], which enforces approximately uniform cluster marginals per mini-batch [38, 39]. For probability vectors $\mathbf{t},\mathbf{a}\in[0,1]^{K}$ with $\sum_{k=1}^{K}t_{k}=\sum_{k=1}^{K}a_{k}=1$, we define the cross-entropy
$\mathrm{CE}(\mathbf{t},\mathbf{a})\;\coloneqq\;-\sum_{k=1}^{K}t_{k}\log a_{k}.$
For notational convenience, we write ${\bm{T}}_{x_{i,j,k,l}}$ and ${\bm{A}}_{x_{i,j,k,l}}$ for the rows of $\bm{T}$} and $\bm{A}$} associated with distorted image $x_{i,j,k,l}$ in Eq. 3. The OT loss has two contributions,
$\mathcal{L}_{\mathrm{ot}}=\frac{1}{N}\sum_{n=1}^{N}\mathrm{CE}\big({\bm{T}}_{n,.},\,{\bm{A}}_{n,.}\big)+\frac{1}{|\mathcal{P}|}\sum_{(u,v)\in\mathcal{P}}\mathrm{CE}\big({\bm{T}}_{u,.},\,{\bm{A}}_{v,.}\big),$ (14)
$\mathcal{P}\;=\;\big\{(x_{i,j,k,l},x_{i,j^{\prime},k,l}):i\in[B],\,k\in[C],\,l\in[L],\,j\neq j^{\prime}\big\}.$
where $N$ is the mini-batch size in Eq. 1. The first term aligns each sample’s assignment to its own balanced target over all $N$ images in the mini-batch. The set $\mathcal{P}$ contains cross-content distorted-image pairs that share the same tiny-batch $i$, composition group $k$, and severity level $l$. The second term averages the corresponding cross-entropy over these pairs.
Soft assignments $\bm{A}$} provide graded membership probabilities, which reduce premature hard assignments and allow clusters to form more flexibly. From $\bm{A}$} we compute an assignment-affinity matrix
${\bm{S}}^{A}_{i,j}\,=\,\sum_{k}{\bm{A}}_{i,k}\,\log{\bm{A}}_{j,k},$
then normalize entries and apply a global sparsity limit to obtain the relation graph
${\bm{G}}^{o}\,=\,F_{K_{g}}\big(\mathrm{norm}({\bm{S}}^{A})\big)\in\mathbb{R}^{N\times N},$ (15)
where $\mathrm{norm}$ maps the matrix entries into $[0,1]$ via global min-max normalization, and $F_{K_{g}}$ keeps the globally largest $K_{g}$ entries, zeroing the rest.
Intuition. OT-based guidance shapes the global topology by encouraging samples with similar prototype memberships to remain close across contents, while soft assignments preserve uncertainty during cluster formation. Prototypes $\bm{C}$} and the encoder $f_{\theta}$ (hence $\bm{H}$}) are updated by gradients of $\mathcal{L}_{\mathrm{ot}}$ together with the overall objective each iteration; $\bm{A}$} and $\bm{T}$} are recomputed from the current features.
III-F Multi-Source Aggregation with Stop-Gradient
Let
$S_{G}\,=\,\{{\bm{G}}^{rd},\,{\bm{G}}^{dd},\,{\bm{G}}^{rr},\,{\bm{G}}^{k},\,{\bm{G}}^{o}\}.$
We apply a stop-gradient to each source and combine them with learned nonnegative coefficients:
$\tilde{{\bm{G}}}^{(i)}$ $=\mathrm{sg}\big({\bm{G}}^{(i)}\big),$ (16)
${\bm{G}}$ $=\sum_{{\bm{G}}^{(i)}\in S_{G}}\omega_{i}\,\tilde{{\bm{G}}}^{(i)},\qquad\omega_{i}\geq 0,\ \sum_{i}\omega_{i}=1,$ (17)
where $\mathrm{sg}(\cdot)$ is the stop-gradient operator (identity in forward, zero gradient in backward). The weights $\omega$ are produced by a lightweight hypernetwork $\Phi$ [40, 19] from simple per-graph statistics,
$\omega$ $=\mathrm{softmax}\Big(\Phi\big(\{\zeta^{(i)}\}_{i=1}^{T}\big)\Big),$ (18)
$\zeta^{(i)}$ $=\big(\mathrm{sum}({\bm{G}}^{(i)}),\ \mathrm{nnz}({\bm{G}}^{(i)})\big),$
where $T=|S_{G}|$, $\{\zeta^{(i)}\}_{i=1}^{T}$ denotes stacking the per-graph statistics over sources, $\mathrm{sum}(\cdot)$ is the total edge weight (connection mass), and $\mathrm{nnz}(\cdot)$ is the number of nonzero entries (sparsity). This hypernetwork aggregates an arbitrary set of sources online, adapting $\omega$ as the graphs evolve during training. The two statistics are inexpensive, scale-agnostic proxies for strength and coverage of each source, providing a practical signal for allocating weight without hand-tuning.
III-G Training Details
We maintain sparse relation graphs via the pruning operators $F_{k_{n}}$ (Top-$k$ per row) and $F_{K_{d}},F_{K_{g}}$ (global Top-$K$). The graph-weighted invariance $\mathcal{L}^{\prime}_{\mathrm{inv}}$ is then evaluated as $\sum_{(i,j):\,{\bm{G}}_{ij}>0}{\bm{G}}_{ij}\,\|{\bm{Z}}_{i,.}-{\bm{Z}}_{j,.}\|_{2}^{2}$, i.e., only over nonzero edges rather than all $N^{2}$ pairs; this respects sparsification and reduces computation. We include the graph regularizer
$\mathcal{R}_{\mathrm{graph}}\;=\;-\sum_{i,j}{\bm{G}}_{i,j}^{\,2},$
to encourage nontrivial mass on the retained edges after sparsification. Under the simplex constraint on $\omega$, this favors informative retained edge weights instead of vanishing graph mass. Our construction choices keep all graph entries in $[0,1]$, which avoids uncontrolled saturation.
III-H End-to-End Optimization and Regression Protocol
The full objective is
$\mathcal{L}_{\mathrm{total}}\,=\,\mathcal{L}_{\mathrm{ssl}}\;+\;\eta\,\mathcal{L}_{\mathrm{ot}}\;+\;\xi\,\mathcal{R}_{\mathrm{graph}}.$ (19)
In each iteration, we construct $\bm{G}$} from the current representations, then update the encoder $f_{\theta}$, projector $g_{\psi}$, the prototype parameters, and the aggregation hypernetwork $\Phi$ by minimizing Eq. 19 while applying stop-gradient through $\bm{G}$}. After pre-training, $g_{\psi}$ is discarded and $f_{\theta}$ is frozen; we then train a linear regressor on the encoder features $\bm{H}$} to predict human opinion scores.
IV Experimental Evaluation
We evaluate SHAMISA on standard NR-IQA benchmarks with a linear-probe protocol [17, 11, 18], report SRCC and PLCC, and study within-dataset performance, cross-dataset transfer, gMAD diagnostics, representation visualizations, ablations, and a full-reference extension.
IV-A Setup
Backbone and pretext
We initialize a ResNet-50 encoder $f_{\theta}$ from ImageNet and pre-train it self-supervised with $\mathcal{L}_{\mathrm{total}}$ (Eq. 19), which couples VICReg-style variance/covariance terms with a graph-weighted invariance term. A 2-layer MLP projector $g_{\psi}$ is used only during pre-training. Unlabeled inputs are degraded online by the compositional distortion engine (Sec. III). Unless otherwise noted, training uses sparse relation graphs ${\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr},{\bm{G}}^{k},{\bm{G}}^{o}$ aggregated by $\Phi$. After pre-training, $g_{\psi}$ is discarded and $f_{\theta}$ is frozen for all evaluations.
Training details
We pre-train for one epoch over the 140k KADIS reference images with SGD. Encoder and projector dimensions are $d_{h}{=}2048$ and $d_{z}{=}256$. Each update uses B=2B{=}2 tiny-batches, R=3R{=}3 references, C=4C{=}4 composition groups, and L=5L{=}5 severity levels, giving $N_{\text{ref}}{=}BR{=}6$, $N_{\text{comp}}{=}CL{=}20$, and $N{=}N_{\text{ref}}(N_{\text{comp}}+1){=}126$ in Eq. 1. Distortion generation uses $M_{d}{=}7$. Graph sparsity and OT settings are $k_{n}{=}31$, $K_{d}{=}4096$, K=32K{=}32, and $\tau_{c}{=}0.1$; for the OT graph, the effective global sparsification budget scales with mini-batch size through $K_{g}{=}8N$. Crop size is $224\times 224$. Remaining details are listed in §A.
IV-B Datasets and Evaluation Protocol
Pre-training data
Following ARNIQA [18], we use the 140k pristine images from KADIS [9] as content sources to enable a fair comparison under a shared pre-training corpus. We generate degradations online with our compositional distortion engine.
Evaluation datasets
We include both synthetic and in-the-wild benchmarks. Synthetic: LIVE [41], CSIQ [42], TID2013 [43], KADID-10K [9]. In-the-wild: FLIVE [29], SPAQ [44]. Brief dataset statistics are available in Appendix D.
Protocol
We follow SSL-IQA practice [17, 11, 18]. For LIVE, CSIQ, TID2013, KADID-10K, and SPAQ, we create 10 train/val/test splits, using reference-disjoint sampling on the synthetic datasets and random 70/10/20 splits on SPAQ. For FLIVE, we use the official split [29]. We train an $\ell_{2}$ ridge regressor [45] on frozen features and select $\alpha_{\mathrm{ridge}}$ from a log-spaced grid of 100 values in $10^{-3},10^{3}]$ by maximizing median validation SRCC across the available splits. At test time, we extract features at full and half scale and concatenate [17], then average predictions over five deterministic crops per image (four corners and center) [10]. We report the median over the repeated splits where applicable. PLCC is computed after the standard four-parameter logistic mapping [46].
IV-C Compared Methods
We compare against traditional NR-IQA (BRISQUE [20], NIQE [21]), codebook models (CORNIA [24], HOSA [25]), and deep supervised NR-IQA (DB-CNN [6], HyperIQA [7], TReS [8], Su et al. [47]). We also include SSL-IQA baselines CONTRIQUE [17], Re-IQA [11], and ARNIQA [18], matching the methods reported in Table I.
IV-D Within-Dataset Performance
TABLE I: Performance comparison on synthetic and in-the-wild NR-IQA datasets. We report SRCC and PLCC. Best and second-best per column are bold and underlined. Prior numbers are taken from their papers or public code where available [17, 11, 18].
Synthetic distortions Authentic distortions
Method Type LIVE CSIQ TID2013 KADID-10K FLIVE SPAQ Average
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
BRISQUE [20] Handcrafted 0.939 0.935 0.746 0.829 0.604 0.694 0.528 0.567 0.288 0.373 0.809 0.817 0.652 0.703
NIQE [21] Handcrafted 0.907 0.901 0.627 0.712 0.315 0.393 0.374 0.428 0.211 0.288 0.700 0.709 0.522 0.572
CORNIA [24] Codebook 0.947 0.950 0.678 0.776 0.678 0.768 0.516 0.558 0.709 0.725
HOSA [25] Codebook 0.946 0.950 0.741 0.823 0.735 0.815 0.618 0.653 0.846 0.852
DB-CNN [6] Supervised 0.968 0.971 0.946 0.959 0.816 0.865 0.851 0.856 0.554 0.652 0.911 0.915 0.841 0.870
HyperIQA [7] Supervised 0.962 0.966 0.923 0.942 0.840 0.858 0.852 0.845 0.535 0.623 0.916 0.919 0.838 0.859
TReS [8] Supervised 0.969 0.968 0.922 0.942 0.863 0.883 0.859 0.858 0.554 0.625
Su et al. [47] Supervised 0.973 0.974 0.935 0.952 0.815 0.859 0.866 0.874
CONTRIQUE [17] SSL+LR 0.960 0.961 0.942 0.955 0.843 0.857 0.934 0.937 0.580 0.641 0.914 0.919 0.862 0.878
Re-IQA [11] SSL+LR 0.970 0.971 0.947 0.960 0.804 0.861 0.872 0.885 0.645 0.733 0.918 0.925 0.859 0.889
ARNIQA [18] SSL+LR 0.966 0.970 0.962 0.973 0.880 0.901 0.908 0.912 0.595 0.671 0.905 0.910 0.869 0.890
SHAMISA (ours) SSL+LR 0.986 0.987 0.981 0.987 0.904 0.919 0.922 0.924 0.610 0.688 0.914 0.920 0.886 0.904
Table I shows that SHAMISA attains the best SRCC and PLCC on LIVE, CSIQ, and TID2013, while remaining competitive on KADID-10K, FLIVE, and SPAQ under the frozen-encoder linear-probe protocol. On KADID-10K, SHAMISA ranks behind CONTRIQUE but ahead of ARNIQA, which is consistent with KADID-10K closely matching the synthetic distortion families emphasized by CONTRIQUE’s pre-training scheme. For authentic images, SHAMISA is second on FLIVE and remains competitive on SPAQ, trailing Re-IQA by a narrow margin. This suggests a stronger balance between synthetic and authentic performance than methods that are markedly stronger on only one side. Overall, SHAMISA achieves the strongest six-dataset average SRCC and PLCC among the SSL methods reported here, with average SRCC/PLCC of 0.886/0.904.
We attribute these gains to two factors: a compositional distortion engine that forms single-factor severity trajectories, and dual-source relation graphs that fuse metadata-driven and structure-intrinsic relations within a graph-weighted VICReg objective. These gains do not come from a materially larger training budget. Under the same KADIS (140k) protocol on an NVIDIA H100 80GB (HBM3), end-to-end pre-training plus evaluation on the six NR-IQA benchmarks takes approximately 12.5 hours for ARNIQA and 13.5 hours for SHAMISA. Re-IQA is more training-intensive, as it additionally trains a content encoder on ImageNet and a separate quality encoder with large augmentation schedules.
IV-E Cross-Dataset Transfer
TABLE II: Cross-dataset SRCC on synthetic NR-IQA benchmarks. Best and second-best per row are bold and underlined. Baseline values are taken from prior reports in ARNIQA and Re-IQA [18, 11].
Training Testing HyperIQA Su et al. CONTRIQUE Re-IQA ARNIQA SHAMISA
LIVE CSIQ 0.744 0.777 0.803 0.795 0.904 0.909
LIVE TID2013 0.541 0.561 0.640 0.588 0.697 0.700
LIVE KADID-10K 0.492 0.506 0.699 0.557 0.764 0.694
CSIQ LIVE 0.926 0.930 0.912 0.919 0.921 0.939
CSIQ TID2013 0.541 0.550 0.570 0.575 0.721 0.729
CSIQ KADID-10K 0.509 0.515 0.696 0.521 0.735 0.711
TID2013 LIVE 0.876 0.892 0.904 0.900 0.869 0.908
TID2013 CSIQ 0.709 0.754 0.811 0.850 0.866 0.870
TID2013 KADID-10K 0.581 0.554 0.640 0.636 0.726 0.779
KADID-10K LIVE 0.908 0.896 0.900 0.892 0.898 0.901
KADID-10K CSIQ 0.809 0.828 0.773 0.855 0.882 0.892
KADID-10K TID2013 0.706 0.687 0.612 0.777 0.760 0.779
Cross-dataset transfer
[…]
Protocol
We follow the standard SSL-IQA linear-probe setting: the encoder is frozen; we train a ridge regressor on the train split of the source dataset, select its regularization on the source val split, and evaluate zero-shot on the target test split with identical multi-scale features and five-crop pooling. We report the median over 10 reference-disjoint splits. This transfer style mirrors practical deployment. A large pool of unlabeled images is used for SSL pre-training, and only a small labeled set from a related source dataset is available to fit a lightweight head. The model must then generalize to a related but distinct target without target labels and without adapting normalization statistics to the target dataset. Full details are provided in Appendix F.
IV-F gMAD Diagnostic on Waterloo
(a) SHAMISA, low bin
(b) SHAMISA, high bin
(c) ARNIQA, low bin
(d) ARNIQA, high bin
Figure 2: gMAD on Waterloo Exploration (distorted pool): panels (a,b) use SHAMISA as the defender and panels (c,d) use ARNIQA as the defender. In each panel, the top and bottom images are the attacker-selected pair from the indicated defender-quality bin, with low-bin cases shown in (a,c) and high-bin cases shown in (b,d). Appendix Table VII reports the corresponding top-1 attacker gaps.
We use the group maximum differentiation (gMAD) protocol [48] as a pairwise diagnostic on the Waterloo Exploration Database [49] using its distorted image pool (no MOS annotations). Following our standard NR-IQA evaluation pipeline, we convert frozen representations into scalar quality predictions by fitting a ridge regressor on KADID-10K and then apply the resulting predictors to all Waterloo images. Given a defender model, we first split the pool into equal-count bins based on the defender’s predicted quality. Within a chosen bin, the attacker then selects the image pair with the largest attacker-side score gap while the defender still scores the pair similarly. Appendix I gives the corresponding protocol details. Fig. 2 shows representative pairs for low and high defender levels in both defender-attacker directions, and Appendix Table VII reports the corresponding top-1 attacker gaps. In this pairwise Waterloo gMAD comparison with KADID-trained ridge heads, attacker gaps are smaller when SHAMISA is the defender than when ARNIQA is the defender, suggesting tighter within-bin quality consistency for SHAMISA in this specific setting.
IV-G t-SNE Representation Visualization
Figure 3: t-SNE visualization of image-level encoder representations on KADID-10K. We extract encoder representations $\mathbf{H}$ (pre-projector) and average across crops to obtain one representation per image, yielding 10,125 image-level points. Points are colored by the coarse distortion family using a shared fixed palette. Compared to ARNIQA, SHAMISA exhibits clearer coarse-family structure for several dominant degradation families, while both models show overlap among visually related families such as brightness and color manipulations. This visualization is a qualitative diagnostic and does not affect the quantitative evaluation protocol.
We visualize the structure of the learned representation space by applying 2D t-SNE [50] to image-level encoder representations $\mathbf{H}$ on KADID-10K [9]. For each image, we extract $\mathbf{H}$ (pre-projector) and average across crops, yielding 10,12510{,}125 points. Fig. 3 colors points by coarse distortion family using a shared, fixed palette for fair visual comparison. Both methods separate several major degradation families, but SHAMISA often forms more coherent regions for prominent categories, whereas ARNIQA shows more inter-family mixing in the central area. We provide additional severity-colored visualizations and subtype-level views (by distortion family) in Appendix G. Across these plots, separability is strongest for families with distinctive artifact statistics (e.g., blur, noise, and compression), while brightness-related manipulations remain more entangled and challenging for both models. Overall, these plots are consistent with SHAMISA’s pre-training encouraging a representation geometry that is coherent across major distortion families while still preserving fine-grained variation within them.
IV-H UMAP Manifold Visualization
(a) ARNIQA
(b) SHAMISA
[…]
Across both models, compositions tend to lie closer to the noise region, consistent with noise being applied after blur and thus remaining visually dominant in the final image. However, the geometry differs markedly between ARNIQA and SHAMISA. The ARNIQA embedding contains a large, diffuse interior region where samples with different blur-to-noise mixtures overlap strongly, yielding a weak and locally inconsistent color gradient. In contrast, SHAMISA appears to produce a more organized manifold, with clearer bands and a smoother progression from blur-dominated (red) to noise-dominated (yellow) compositions, as well as fewer ambiguous mixed-mixture points concentrated in the middle. This qualitative diagnostic suggests that SHAMISA more explicitly parameterizes the two distortion axes under composition: blur and noise vary more smoothly, and with less central mixing than ARNIQA.
IV-I Ablations
TABLE III: Main ablations. Median over 10 splits per dataset; then averaged. SRCCs/PLCCs average over synthetic $\{\mathrm{LIVE},\mathrm{CSIQ},\mathrm{TID2013},\mathrm{KADID\mbox{-}10K}\}$, SRCCa/PLCCa over authentic $\{\mathrm{FLIVE},\mathrm{SPAQ}\}$, and SRCCg/PLCCg denote the average over the six per-dataset median SRCC/PLCC values. Full ablation catalog is reported in Appendix Table VIII.
Variant SRCCg PLCCg SRCCa PLCCa SRCCs PLCCs
A0: SHAMISA 0.8862 0.9042 0.7620 0.8040 0.9483 0.9543
(A) Aggregation & regularization
A1: Binary $\bm{G}$} 0.8767 0.8924 0.7475 0.7821 0.9374 0.9457
A3: No $\Phi$ (fixed linear mixture) 0.8759 0.8951 0.7479 0.7896 0.9361 0.9457
(B) Graph sources
B1: No ${\bm{G}}^{o}$ (OT graph) 0.8600 0.8787 0.7661 0.8069 0.8995 0.9092
B7: Metadata only $\{{\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr}\}$ 0.8467 0.8648 0.7634 0.8010 0.8801 0.8905
B6: ${\bm{G}}^{k}$ only 0.7891 0.8018 0.7577 0.7825 0.7922 0.8017
B8: Structure only $\{{\bm{G}}^{k},{\bm{G}}^{o}\}$ 0.8753 0.8961 0.7419 0.7891 0.9387 0.9475
(D) Distortion engine
D1: $M_{d}{=}1$ 0.8775 0.9077 0.7259 0.7684 0.9519 0.9783
D2: Discrete severities (L=5L{=}5) 0.8767 0.8987 0.7456 0.7890 0.9386 0.9518
D4: Fixed order $\pi$ 0.8587 0.8712 0.7434 0.7668 0.9113 0.9215
(E) Objective diagnostic
E1: No $\mathcal{L}_{\mathrm{ot}}$ 0.7887 0.8009 0.7597 0.7836 0.7904 0.7996
Ablation design
We ablate SHAMISA along complementary axes that map directly to our main design choices: (A) how we aggregate multi-source relations into the soft matrix $\bm{G}$} (learned end-to-end mixing $\Phi$ vs fixed mixing; regularization), (B) which relation sources are necessary (structure graphs ${\bm{G}}^{k},{\bm{G}}^{o}$ vs metadata graphs ${\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr}$), (C) how we sparsify the global OT graph (global vs row-wise truncation), (D) how the compositional distortion engine defines the self-supervised training distribution (composition complexity, severity modeling, ordering), and (E–F) objective diagnostics that isolate the contribution of OT alignment and severity-dependent edge weighting. Each variant changes only minimal factors relative to the reference configuration (A0). We evaluate on six NR-IQA datasets (LIVE, CSIQ, TID2013, KADID-10K, FLIVE, SPAQ) with 10 random splits per dataset, reporting the median test SRCC/PLCC per dataset, then summarizing synthetic and authentic groups separately together with a global comparison score. For readability, Table III highlights the most diagnostic variants; the complete catalog is deferred to Appendix Table VIII.
Results and analysis
[…]
Finally, ablations of relation aggregation indicate that how relations are mixed matters, but the effects are necessarily smaller than removing relations entirely. Binarizing $\bm{G}$} (A1) slightly reduces SRCCg (0.8767), supporting soft relation strengths for graded supervision. Replacing learned mixing $\Phi$ with fixed linear mixing (A3) decreases SRCCg to 0.8759, consistent with $\Phi$ learning to reweight heterogeneous relation sources. Removing the graph regularizer $\mathcal{R}_{\mathrm{graph}}$ (A2) yields only a small drop in SRCCg (0.8813), suggesting that it mainly improves stability and calibration rather than acting as a dominant source of the gains.
IV-J Pre-training Dynamics
We track downstream NR-IQA performance during SSL pre-training by periodically freezing the encoder and fitting a lightweight regressor on each target dataset. Frozen checkpoints are evaluated every 5000 training steps. Fig. 5 shows that SRCC improves rapidly in the early phase and then saturates across the representative benchmark datasets shown here, indicating that the learned representation becomes useful for quality prediction relatively early in pre-training.
Importantly, the best checkpoint can be dataset-dependent: while some datasets continue to improve or remain stable as pre-training proceeds, others exhibit a mild peak-and-decline behavior, which is expected because the SSL objective is not optimized directly for SRCC on any single benchmark. As a deployment note, when a small labeled validation set is available in the target domain, it can be used to select a checkpoint for that domain. Unless stated otherwise, all results reported in Tables I and II use the fixed A0 checkpoint; this section is a diagnostic and is not used for model selection. Additional diagnostic figures are provided in Appendix K.
Figure 5: SRCC vs. training steps during SSL pre-training (mean with $\pm$ std error bars over 10 splits) across representative NR-IQA benchmarks. Curves typically rise early and then saturate, while the best checkpoint can vary by dataset since the SSL objective is not identical to downstream SRCC.
IV-K Hyperparameter Sensitivity
Figure 6: One-at-a-time sweeps around the SHAMISA reference A0, reporting SRCC across six NR-IQA benchmarks in a single merged view. Columns correspond to $(d_{h},d_{z})$, $(B,\xi)$, $(\alpha,\beta)$, $(\gamma,\eta)$, and $(k_{n},K)$, with a shared legend across all panels.
We analyze hyperparameter sensitivity by varying one knob at a time around our A0 configuration while keeping the rest fixed. Fig. 6 shows that most capacity and graph-construction choices admit broad plateaus: $d_{h}$, $d_{z}$, and $k_{n}$ typically yield stable SRCC across a wide range, and increasing $K$ mainly saturates after moderate values. In contrast, the OT alignment coefficient $\eta$ is the most influential knob across datasets, with the largest SRCC swings on synthetic distortion benchmarks such as TID2013 and KADID-10K, suggesting that overly weak alignment under-utilizes relational guidance, whereas overly strong alignment can over-constrain representations. Among VICReg terms, the invariance weight $\gamma$ exhibits the clearest sensitivity, particularly on FLIVE, whereas the variance and covariance weights $\alpha$ and $\beta$ are comparatively robust across the tested ranges. Overall, these sweeps indicate that SHAMISA is robust across broad ranges of capacity and graph-design choices, while a smaller set of alignment-related coefficients, especially $\eta$ and $\gamma$, remains important for fine control. Appendix L reports the corresponding SRCC sensitivity summary.
FR extension
We also propose a full-reference variant, SHAMISA-FR, that reuses the frozen no-reference encoder and uses the absolute feature difference $|h_{\mathrm{ref}}{-}h_{\mathrm{dist}}|$ between pristine and distorted images as the regressor input. This variant requires no additional backbone pre-training and attains competitive FR-IQA accuracy; details are provided in Appendix E, with the main comparison in Appendix Table VI.
Reproducibility
Additional implementation details, optimization settings, evaluation protocols, and supplementary diagnostics are provided in Appendices A, F, K, and L. The project repository is available at https://github.com/Mahdi-Naseri/SHAMISA/.
V Conclusion
We presented SHAMISA, a non-contrastive self-supervised framework for NR-IQA that learns unified distortion-aware and content-aware representations. SHAMISA couples a compositional distortion engine with single-factor variation to generate controllable severity trajectories, and supervises learning through dual-source relation graphs that fuse Metadata-Driven Graphs with Structurally Intrinsic Graphs. These relations drive a graph-weighted VICReg-style invariance with clustering-guided supervision. In each iteration, we construct the graphs from the current representations and then update the model with stop-gradient. This yields a frozen encoder whose features support accurate quality prediction with a linear regressor. Across synthetic and authentic benchmarks, SHAMISA achieves the strongest overall average among the SSL methods compared here, with improved cross-dataset generalization and stronger gMAD diagnostics, while using no human quality labels during pre-training. A full-reference variant, SHAMISA-FR, reuses the same frozen encoder without additional backbone training and attains competitive FR-IQA accuracy. Future work includes exploiting the learned topology for blind image restoration and extending the framework to video quality assessment.
Supplementary Material for “SHAMISA: SHAped Modeling of Implicit Structural Associations for Self-supervised No-Reference Image Quality Assessment”
Mahdi Naseri and Zhou Wang
Appendix A Architectures and Optimization
Table IV summarizes the fixed SHAMISA pre-training configuration used across all experiments; unless stated otherwise, downstream evaluation selects the ridge regularization $\alpha_{\mathrm{ridge}}$ per dataset on the validation split, as described in Sec. IV-B.
TABLE IV: Essential final hyperparameters used for SHAMISA pre-training. A single configuration is shared across all datasets (no dataset-specific tuning during SSL).
Group Hyperparameters
Architecture ResNet-50 encoder; $d_{h}=2048$; projector output $d_{z}=256$.
Optimization SGD with $\mathrm{lr}=1.5\times 10^{-3}$, momentum $0.9$, weight decay $10^{-4}$, and cosine annealing warm restarts.
Loss VICReg-style weights $(\alpha,\beta,\gamma)=(11.98,57.21,88.37)$; OT alignment weight $\eta=0.4906$; graph regularizer weight $\xi=0.0342$.
Graphs kNN graph: $k_{n}=31$; metadata graph sparsifier: $K_{d}=4096$; ref-ref edge weight $w_{rr}=0.5766$; OT graph: $K=32$, $\tau_{c}=0.1$; OT global sparsifier: $K_{g}=8N$.
Distortion engine Per update: $B=2$ tiny-batches, $R=3$ references, $C=4$ composition groups, $L=5$ severity levels; crop size $224$; max distortions per sample $M_{d}=7$.
Implementation detail for graph construction: diagonal self-edges are removed from all source graphs before aggregation. We therefore do not use self-invariance edges in Eq. 7.
Appendix B Compositional Distortion Engine
We use the compositional engine of Sec. III-B with $M_{d}=7$ and per-update structure $B=2$ tiny-batches, each with $R=3$ pristine references and $C=4$ distortion composition groups. Each group forms a single-factor trajectory with $L=5$ levels, generating images $x_{i,j,k,l}$ as defined in Eq. 3. Normalized severities are sampled i.i.d. by drawing $\epsilon\sim\mathcal{N}(0,0.5^{2})$ and setting $s=\min(1,|\epsilon|)$; within each group, one distortion coordinate varies across $L$ i.i.d. draws $\{s^{(l)}\}_{l=1}^{L}$, which are kept in sampled order and are not forced to include either endpoint of the severity range, while all other coordinates are held at their baseline sampled values. Distortion functions are applied sequentially in a uniformly random order (Eq. 2).
Appendix C Evaluation Details
Following prior SSL-IQA evaluation practice [17, 11, 18], we use two scales per image (full and half) and five deterministic crops per scale [17, 10], with ridge regularization selected on the validation split as in Sec. IV-B. We report the median over 10 random splits where applicable. On synthetic datasets, the splits are reference-disjoint following [18, 17, 11]. For FLIVE we use the official split [29], and for SPAQ we use random 70/10/20 splits. PLCC follows the same four-parameter logistic mapping specified in Sec. IV-B.
Appendix D Datasets
We evaluate on four synthetic IQA benchmarks and two authentic in-the-wild IQA benchmarks. Table V summarizes the dataset sizes, label types, and split notes used throughout the paper.
Among the authentic datasets, FLIVE provides large-scale in-the-wild images with both image-level and patch-level quality annotations [29], whereas SPAQ focuses on smartphone photography captured under diverse real-world conditions [44]. For SPAQ, we resize the shorter side to 512 as in [44].
TABLE V: Dataset statistics for the NR-IQA benchmarks. Synthetic datasets are reported by reference count and distortion inventory, whereas authentic datasets are summarized by image count and label type.
Dataset Type Pristine refs Distortion types Images Label type Split note
LIVE [41] Synthetic 29 5 779 DMOS Reference-disjoint
CSIQ [42] Synthetic 30 6 866 DMOS Reference-disjoint
TID2013 [43] Synthetic 25 24 3000 MOS Reference-disjoint
KADID-10K [9] Synthetic 81 25 10,125 MOS Reference-disjoint
FLIVE [29] In-the-wild 39,808 MOS + patch scores Official split
SPAQ [44] In-the-wild 11,125 MOS Random 70/10/20
Appendix E SHAMISA-FR: Full-Reference Variant
TABLE VI: Full-reference IQA comparison on synthetic distortion datasets. Reported rows are taken from prior work. SHAMISA-FR uses the same fixed-encoder evaluation protocol as our NR setting, and each value is the median over 10 random splits. Best and second-best values in each column are shown in bold and underlined.
Method Type LIVE CSIQ TID2013 KADID-10K
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
PSNR Traditional 0.881 0.868 0.820 0.824 0.643 0.675 0.677 0.680
SSIM [52] Traditional 0.921 0.911 0.854 0.835 0.642 0.698 0.641 0.633
FSIM [53] Traditional 0.964 0.954 0.934 0.919 0.852 0.875 0.854 0.850
VSI [54] Traditional 0.951 0.940 0.944 0.929 0.902 0.903 0.880 0.878
PieAPP [55] Deep learning 0.915 0.905 0.900 0.881 0.877 0.850 0.869 0.869
LPIPS [36] Deep learning 0.932 0.936 0.884 0.906 0.673 0.756 0.721 0.713
DISTS [56] Deep learning 0.953 0.954 0.942 0.942 0.853 0.873
DRF-IQA [57] Deep learning 0.983 0.983 0.964 0.960 0.944 0.942
CONTRIQUE-FR [17] SSL + LR 0.966 0.966 0.956 0.964 0.909 0.915 0.946 0.947
Re-IQA-FR [58] SSL + LR 0.969 0.974 0.961 0.962 0.920 0.921 0.933 0.936
ARNIQA-FR [18] SSL + LR 0.969 0.972 0.971 0.975 0.898 0.901 0.920 0.919
SHAMISA-FR SSL + LR 0.962 0.970 0.968 0.982 0.909 0.906 0.899 0.892
Setup
SHAMISA-FR reuses the pre-trained $f_{\theta}$ from the no-reference setting and keeps it frozen. A key point is that the FR variant adds only a lightweight regressor on top of frozen features, so it requires no additional backbone pre-training and only minimal extra computation. Given $(x_{\mathrm{ref}},x_{\mathrm{dist}})$, we compute
$h_{\mathrm{ref}}=f_{\theta}(x_{\mathrm{ref}}),\qquad h_{\mathrm{dist}}=f_{\theta}(x_{\mathrm{dist}}),$
then form $u=\lvert h_{\mathrm{ref}}-h_{\mathrm{dist}}\rvert$. A linear regressor is trained on $u$ to predict the quality label (DMOS for LIVE/CSIQ, MOS for TID2013/KADID-10K) without updating $f_{\theta}$.
Protocol
We evaluate on four synthetic FR-IQA datasets (LIVE, CSIQ, TID2013, KADID-10K), using the same split logic as in our NR evaluation, and report median SRCC/PLCC over 10 random splits. For context, Table VI also lists representative FR-IQA baselines as reported in prior work [17, 58, 18].
Results
Table VI shows that SHAMISA-FR is most competitive on CSIQ among the SSL-based FR rows, while remaining competitive on LIVE and trailing more clearly on TID2013 and KADID-10K, which cover a broader mix of distortions. Supervised FR baselines remain stronger because they optimize full networks directly on the target FR datasets, whereas SHAMISA-FR deliberately tests how far a frozen no-reference representation can transfer with only a lightweight head. As the backbone is fixed, these results primarily reflect the pre-trained representation.
Appendix F Cross-Dataset Evaluation Protocol
Practical motivation
The protocol is designed to reflect common deployment constraints in no-reference quality assessment. In practice, one has abundant unlabeled images for self-supervised pre-training, a limited labeled dataset from a related source domain to fit a lightweight head, and a target distribution that is similar but not identical, with little to no supervision. Freezing the encoder, avoiding any target-specific normalization, and disallowing target-time hyperparameter tuning isolate representation quality and prevent leakage across domains.
Setting
We assess zero-shot transfer across synthetic NR-IQA datasets {LIVE, CSIQ, TID2013, KADID-10K}. The encoder $f_{\theta}$ is frozen for all evaluations and pre-trained once under the setup of Sec. IV-A. Only a ridge regressor is trained per source dataset.
Source-target procedure
For each source dataset $D_{s}$, we create 10 reference-disjoint 70/10/20 train/val/test splits on $D_{s}$. We extract frozen features at two scales and five crops per scale, concatenate features, and train a ridge regressor on the train split of $D_{s}$. The regularization coefficient is selected once on the val split of $D_{s}$ via a logarithmic grid from $10^{-3}$ to $10^{3}$. The selected regressor is then evaluated, without any modification, on the test split of every target $D_{t}\neq D_{s}$.
Preprocessing and normalization
All synthetic datasets in this protocol use the same resizing, cropping, and color normalization pipelines. No statistics from $D_{t}$ are computed or applied. In particular, we do not apply target-specific normalization or retune hyperparameters.
Metrics and aggregation
We report SRCC on raw predictions. For each $D_{s}\to D_{t}$, the headline result is the median over the 10 source splits.
Contrast to some prior practice
Some SSL-IQA works may emphasize within-dataset or mixed-dataset training. Our protocol enforces strict zero-shot transfer $D_{s}\to D_{t}$ with a fixed encoder and hyperparameters chosen on $D_{s}$, isolating representation quality rather than target adaptation.
Appendix G Additional t-SNE Visualizations
Protocol
We compute t-SNE [50] on KADID-10K [9] using image-level encoder representations $\mathbf{H}$ (pre-projector). For each image, we extract features from multiple crops and average them, resulting in 10,125 image-level representations. Before t-SNE, we apply scikit-learn PCA to 50 dimensions with ‘random_state=123‘ and ‘svd_solver=auto‘. We then run 2D t-SNE using scikit-learn TSNE with a fixed seed shared across models: ‘random_state=123‘, ‘perplexity=30‘, ‘learning_rate=200‘, ‘n_iter=2000‘, and ‘init=pca‘, with default ‘metric=euclidean‘ and ‘early_exaggeration=12.0‘. We apply no additional feature normalization before PCA or t-SNE. Unless otherwise noted, the plotted images and t-SNE settings are identical across models; only the extracted representations differ. All visualizations are qualitative diagnostics and are not used for model selection.
Additional visualizations
Fig. 7 colors the global KADID-10K embedding by severity level and shows that several coarse families exhibit smooth local severity progressions in representation space, even though visually related distortions still overlap. Fig. 8 then focuses on SHAMISA at the subtype level. Blur, compression, and spatial distortions show the clearest subtype-level separation, whereas color and brightness-related manipulations remain more entangled.
Figure 7: t-SNE colored by distortion severity on KADID-10K. Colors indicate severity levels 1 to 5 for ARNIQA (left) and SHAMISA (right).
Figure 8: Subtype-colored t-SNE views for SHAMISA, including six KADID-10K distortion families.
Appendix H UMAP Protocol Details
[…]
Appendix I gMAD Protocol Details
For each model, we compute scalar quality predictions on the full Waterloo Exploration distorted pool using a frozen encoder and a ridge regressor trained on KADID-10K, following the feature extraction and normalization in Sec. IV-B. For a given defender $f_{\mathrm{def}}$ and attacker $f_{\mathrm{att}}$, we bin all images into $N_{\mathrm{bins}}{=}10$ equal-count levels based on $f_{\mathrm{def}}$ predictions and select the lowest and highest bins. Within each selected bin, we select the top-1 attacker-gap pair, defined as the pair with maximal attacker score separation while remaining in the same defender-defined level, and report it for visualization. All reported figures use the Waterloo distorted subset. Table VII summarizes the corresponding top-1 attacker gaps for the four defender/quality-level cases.
TABLE VII: gMAD summary on Waterloo Exploration (distorted pool) between SHAMISA and ARNIQA. Attacker gap denotes the attacker-side score separation of the selected pair within the defender-defined bin; lower is better.
Defender, Quality Level Bin size Top-1 attacker gap
ARNIQA, high 9488 8.2717
ARNIQA, low 9488 8.6782
SHAMISA, high 9488 4.1860
SHAMISA, low 9488 3.1356
Appendix J Additional Ablations
Table VIII reports the complete ablation catalog, including finer-grained graph removals (B2–B5), OT sparsification (C1), the additional distortion-engine setting D3, and severity-dependent edge weighting (F1). Two patterns are worth noting. (1) Single metadata edge removals (B3–B5) each cause only small changes, while metadata-only (B7) remains substantially below A0, indicating metadata relations are individually weak but collectively useful through redundancy and coverage of diverse distortion semantics. (2) Severity-dependent edge weighting is helpful but secondary: replacing $\phi,\hat{\phi}$ with the identity (F1) reduces SRCCg to 0.8736, whereas removing OT alignment (E1) drops to 0.7887, reinforcing that OT-guided global structure is the central mechanism while severity shaping refines it.
TABLE VIII: Full ablation catalog. For each dataset we run 10 random splits and report the median test SRCC/PLCC across splits, then average within groups: synthetic $\{\mathrm{LIVE},\mathrm{CSIQ},\mathrm{TID2013},\mathrm{KADID\mbox{-}10K}\}$, authentic $\{\mathrm{FLIVE},\mathrm{SPAQ}\}$, and SRCCg/PLCCg denote the average over the six per-dataset median SRCC/PLCC values.
Variant SRCCg PLCCg SRCCa PLCCa SRCCs PLCCs
A0: SHAMISA 0.8862 0.9042 0.7620 0.8040 0.9483 0.9543
(A) Graph weighting & regularization
A1: Binary $\bm{G}$} 0.8767 0.8924 0.7475 0.7821 0.9374 0.9457
A2: No $\mathcal{R}_{\mathrm{graph}}$ 0.8813 0.8981 0.7471 0.7908 0.9450 0.9497
A3: No $\Phi$ (fixed linear mixture) 0.8759 0.8951 0.7479 0.7896 0.9361 0.9457
A4: No stop-grad through $\Phi$ 0.8797 0.8980 0.7521 0.7950 0.9395 0.9469
(B) Source graph contributions
B1: No ${\bm{G}}^{o}$ (OT graph) 0.8600 0.8787 0.7661 0.8069 0.8995 0.9092
B2: No ${\bm{G}}^{k}$ (kNN graph) 0.8815 0.8989 0.7475 0.7890 0.9452 0.9520
B3: No ${\bm{G}}^{rd}$ (ref–dist metadata) 0.8765 0.8931 0.7484 0.7865 0.9366 0.9443
B4: No ${\bm{G}}^{dd}$ (dist–dist metadata) 0.8771 0.8937 0.7494 0.7867 0.9370 0.9451
B5: No ${\bm{G}}^{rr}$ (ref–ref metadata) 0.8814 0.8976 0.7462 0.7854 0.9458 0.9520
B6: ${\bm{G}}^{k}$ only 0.7891 0.8018 0.7577 0.7825 0.7922 0.8017
B7: Metadata only $\{{\bm{G}}^{rd},{\bm{G}}^{dd},{\bm{G}}^{rr}\}$ 0.8467 0.8648 0.7634 0.8010 0.8801 0.8905
B8: Structure only $\{{\bm{G}}^{k},{\bm{G}}^{o}\}$ 0.8753 0.8961 0.7419 0.7891 0.9387 0.9475
(C) Sparsity operators & topology shaping
C1: Row-wise Top-$k$ in ${\bm{G}}^{o}$ (k=8k{=}8) 0.8734 0.8898 0.7500 0.7918 0.9307 0.9359
(D) Distortion engine
D1: $M_{d}{=}1$ 0.8775 0.9077 0.7259 0.7684 0.9519 0.9783
D2: Discrete severities (L=5L{=}5) 0.8767 0.8987 0.7456 0.7890 0.9386 0.9518
D3: Uniform severity sampling 0.8743 0.8955 0.7439 0.7946 0.9359 0.9432
D4: Fixed composition order $\pi$ 0.8587 0.8712 0.7434 0.7668 0.9113 0.9215
(E) Objective form & diagnostics
E1: No $\mathcal{L}_{\mathrm{ot}}$ 0.7887 0.8009 0.7597 0.7836 0.7904 0.7996
(F) Severity mapping diagnostics
F1: Identity $\phi,\hat{\phi}$ 0.8736 0.8937 0.7459 0.7914 0.9335 0.9423
Appendix K Additional Pre-Training Dynamics
K-A Downstream SRCC During SSL Pre-Training
We track downstream SRCC of the frozen pre-trained encoder across pre-training snapshots using the same lightweight regressor as in the main experiments. Fig. 5 complements the main text by showing that SRCC improves rapidly in the early phase and then largely saturates across datasets, while the final checkpoint is not uniformly optimal for every dataset. This figure serves as a compact convergence record rather than a separate endpoint-results table.
K-B Encoder Representations and Projector Embeddings During SSL Pre-Training
We additionally track correlation and invariance for both encoder representations $\bm{H}$} and projector embeddings $\bm{Z}$}, together with rank for $\bm{H}$}, during SSL pre-training. Fig. 9 shows that the correlation proxy drops quickly and then stabilizes for both $\bm{H}$} and $\bm{Z}$}, consistent with progressively reduced feature redundancy. Fig. 10 shows that the effective dimensionality of $\bm{H}$} evolves during training and then settles after the early phase, while Fig. 11 shows a similar decrease and stabilization in the invariance proxy.
(a) corr for $\bm{H}$} vs. steps.
(b) corr for $\bm{Z}$} vs. steps.
Figure 9: Feature decorrelation proxy (corr) over SSL pre-training. Both $\bm{H}$} and $\bm{Z}$} show a rapid early drop followed by stabilization, consistent with progressively reduced redundancy across feature dimensions.
Figure 10: Effective dimensionality proxy (rank) for $\bm{H}$} over SSL pre-training. The encoder representation uses a substantial number of active dimensions throughout training and then stabilizes after the early phase.
(a) inv for $\bm{H}$} vs. steps.
(b) inv for $\bm{Z}$} vs. steps.
Figure 11: Invariance proxy (inv) over SSL pre-training. Curves generally decrease and then stabilize, consistent with paired views becoming closer in representation space as training proceeds.
Appendix L Hyperparameter Sensitivity Details
We perform one-at-a-time sweeps around the A0 reference configuration, varying a single hyperparameter while keeping all others fixed. Table IX summarizes SRCC sensitivity on LIVE, CSIQ, TID2013, KADID-10K, FLIVE, and SPAQ. For each hyperparameter, $\Delta$ is the largest minus smallest SRCC across the sweep. Larger $\Delta$ indicates higher sensitivity.
TABLE IX: Sensitivity summary for SRCC. For each hyperparameter, $\Delta$ is the largest minus smallest SRCC across the sweep. Larger $\Delta$ indicates higher sensitivity. The most sensitive and second most sensitive hyperparameters per dataset are shown in bold and underlined.
HP LIVE CSIQ TID2013 KADID-10K FLIVE SPAQ Mean
$d_{h}$ 0.012 0.003 0.003 0.006 0.004 0.007 0.006
$d_{z}$ 0.006 0.002 0.002 0.000 0.003 0.007 0.003
$B$ 0.002 0.001 0.003 0.005 0.001 0.001 0.002
$k_{n}$ 0.007 0.009 0.006 0.016 0.007 0.001 0.008
$K$ 0.019 0.035 0.061 0.030 0.007 0.005 0.026
$\gamma$ 0.017 0.045 0.089 0.070 0.033 0.001 0.042
$\alpha$ 0.003 0.004 0.006 0.010 0.003 0.001 0.004
$\beta$ 0.005 0.004 0.006 0.005 0.007 0.003 0.005
$\xi$ 0.010 0.002 0.003 0.013 0.000 0.000 0.005
$\eta$ 0.054 0.096 0.208 0.226 0.035 0.001 0.104

### 17 Mar 2026 v2

@@
 From this, we build two categories of relation graphs: (i) Metadata-Driven Graphs, which encode pairwise similarity based on distortion metadata, encouraging images with similar degradations to lie close in the learned manifold while inducing controlled representational shifts as distortion severity varies; and (ii) Structurally Intrinsic Graphs, constructed from the latent feature space using $k$-nearest neighbors (kNN) and deep clustering.
-These graphs supervise a non-contrastive VICReg-style objective, optimized in a single update that constructs relation graphs from the current representations and applies stop-gradient through graph construction.
+These graphs supervise a non-contrastive VICReg-style objective with graph-weighted invariance.
+With stop-gradient applied to graph construction, each iteration builds relation graphs from the current representations and updates the model under the resulting objective in a single optimization step.
 SHAMISA unifies content-dependent and distortion-dependent learning within a single relational framework, generalizing the rigid pairing schemes used in prior SSL-IQA methods.
@@
 Crop size is $224\times 224$.
-Remaining details are listed in §A.
+Remaining details are listed in Supplementary Appendix A.
 IV-B Datasets and Evaluation Protocol
@@
 In-the-wild: FLIVE [29], SPAQ [44].
-Brief dataset statistics are available in Appendix D.
+Brief dataset statistics are available in Supplementary Appendix D.
 Protocol
@@
 The model must then generalize to a related but distinct target without target labels and without adapting normalization statistics to the target dataset.
-Full details are provided in Appendix F.
+Full details are provided in Supplementary Appendix F.
 IV-F gMAD Diagnostic on Waterloo
@@
 In each panel, the top and bottom images are the attacker-selected pair from the indicated defender-quality bin, with low-bin cases shown in (a,c) and high-bin cases shown in (b,d).
-Appendix Table VII reports the corresponding top-1 attacker gaps.
+Supplementary Table VII reports the corresponding top-1 attacker gaps.
 We use the group maximum differentiation (gMAD) protocol [48] as a pairwise diagnostic on the Waterloo Exploration Database [49] using its distorted image pool (no MOS annotations).
@@
 Within a chosen bin, the attacker then selects the image pair with the largest attacker-side score gap while the defender still scores the pair similarly.
-Appendix I gives the corresponding protocol details.
+Supplementary Appendix I gives the corresponding protocol details.
 Fig.
-2 shows representative pairs for low and high defender levels in both defender-attacker directions, and Appendix Table VII reports the corresponding top-1 attacker gaps.
+2 shows representative pairs for low and high defender levels in both defender-attacker directions, and Supplementary Table VII reports the corresponding top-1 attacker gaps.
 In this pairwise Waterloo gMAD comparison with KADID-trained ridge heads, attacker gaps are smaller when SHAMISA is the defender than when ARNIQA is the defender, suggesting tighter within-bin quality consistency for SHAMISA in this specific setting.
@@
 Both methods separate several major degradation families, but SHAMISA often forms more coherent regions for prominent categories, whereas ARNIQA shows more inter-family mixing in the central area.
-We provide additional severity-colored visualizations and subtype-level views (by distortion family) in Appendix G.
+We provide additional severity-colored visualizations and subtype-level views (by distortion family) in Supplementary Appendix G.
 Across these plots, separability is strongest for families with distinctive artifact statistics (e.g., blur, noise, and compression), while brightness-related manipulations remain more entangled and challenging for both models.
@@
 SRCCs/PLCCs average over synthetic $\{\mathrm{LIVE},\mathrm{CSIQ},\mathrm{TID2013},\mathrm{KADID\mbox{-}10K}\}$, SRCCa/PLCCa over authentic $\{\mathrm{FLIVE},\mathrm{SPAQ}\}$, and SRCCg/PLCCg denote the average over the six per-dataset median SRCC/PLCC values.
-Full ablation catalog is reported in Appendix Table VIII.
+Full ablation catalog is reported in Supplementary Table VIII.
 Variant SRCCg PLCCg SRCCa PLCCa SRCCs PLCCs
@@
 We evaluate on six NR-IQA datasets (LIVE, CSIQ, TID2013, KADID-10K, FLIVE, SPAQ) with 10 random splits per dataset, reporting the median test SRCC/PLCC per dataset, then summarizing synthetic and authentic groups separately together with a global comparison score.
-For readability, Table III highlights the most diagnostic variants; the complete catalog is deferred to Appendix Table VIII.
+For readability, Table III highlights the most diagnostic variants; the complete catalog is deferred to Supplementary Table VIII.
 Results and analysis
@@
 Unless stated otherwise, all results reported in Tables I and II use the fixed A0 checkpoint; this section is a diagnostic and is not used for model selection.
-Additional diagnostic figures are provided in Appendix K.
+Additional diagnostic figures are provided in Supplementary Appendix K.
 Figure 5: SRCC vs.
@@
 Overall, these sweeps indicate that SHAMISA is robust across broad ranges of capacity and graph-design choices, while a smaller set of alignment-related coefficients, especially $\eta$ and $\gamma$, remains important for fine control.
-Appendix L reports the corresponding SRCC sensitivity summary.
+Supplementary Appendix L reports the corresponding SRCC sensitivity summary.
 FR extension
 We also propose a full-reference variant, SHAMISA-FR, that reuses the frozen no-reference encoder and uses the absolute feature difference $|h_{\mathrm{ref}}{-}h_{\mathrm{dist}}|$ between pristine and distorted images as the regressor input.
-This variant requires no additional backbone pre-training and attains competitive FR-IQA accuracy; details are provided in Appendix E, with the main comparison in Appendix Table VI.
+This variant requires no additional backbone pre-training and attains competitive FR-IQA accuracy; details are provided in Supplementary Appendix E, with the main comparison in Supplementary Table VI.
 Reproducibility
-Additional implementation details, optimization settings, evaluation protocols, and supplementary diagnostics are provided in Appendices A, F, K, and L.
+Additional implementation details, optimization settings, evaluation protocols, and supplementary diagnostics are provided in Supplementary Appendix A, Supplementary Appendix F, Supplementary Appendix K, and Supplementary Appendix L.
 The project repository is available at https://github.com/Mahdi-Naseri/SHAMISA/.